OpenStudy (anonymous):

: PDE's: / : Waves: A string has its ends fixed at x=0 and x=L. It is displaced a distance h at its midpoint and then released. Formulate a boundary value problem for the displacement y(x,t) of any point x of the string at time t.

OpenStudy (anonymous):

dont reckon anyone is online that are capable, source: asked domain ? earlier wish could help

OpenStudy (anonymous):

Here's what I got: Let y = y=(x,t) and a^2=tau/mu. Then y_tt = a^2(y_xx) and y(L/2,0)=h, y_t(x,0)=0, y(0,t)=0, y(L,t)=0. Please confirm if this is correct or not.

OpenStudy (anonymous):

@tmauldin: haha thanks, i appreciate it. that's alright, my homework isn't due until 11/21 anyway, I can wait or repost this tomorrow or sometime again so that someone online who's capable could see it.